OpenStudy (anonymous):

Rewrite in vertex form: y=x^2+14x+50 Please help me understand.

OpenStudy (anonymous):

hi, still need a hand?

OpenStudy (anonymous):

to get the eqution into vertex form, we will need to 'complete the square' do you know how to do this?

OpenStudy (anonymous):

No sir I do not lol, I'm just really messed up on this.

OpenStudy (anonymous):

no problem, we can get you all set up. here is what we are going to do, we need to 'rewrite' this: y=x^2+14x+50 which has a 'form' of y=Ax^2+Bx+c and we need to turn it into a 'form' of y=a(x-h)^2 + k do these 'forms' look familiar?

OpenStudy (anonymous):

Yes they do, I seen that method before

OpenStudy (anonymous):

ok then to rewrite the equation into the 'vertex form' we will 'complete the square'

OpenStudy (anonymous):

we take the coefficient that is infront of the 'x' term, and divide it by 2